OpenStudy (anonymous):

find value of x in the interval [0,2π/3] such that 2cos(3x)-1=0 . I got up to: cos(3x) =1/2 but after that I am lost :(

OpenStudy (anonymous):

so you know for which x you get \[\cos(x)=\frac{1}{2}\]?

OpenStudy (anonymous):

i believe on value sis \[\frac{\pi}{3}\] so that means \[3x=\frac{\pi}{3}\] solve to get \[x=\frac{\pi}{9}\] as one answer

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

first you need a value that gives \[\cos(x)=\frac{1}{2}\] right?

OpenStudy (anonymous):

then replace x by 3x and solve. if you know that \[\cos(\frac{\pi}{3})=\frac{1}{2}\] that means \[3x=\frac{\pi}{3}\] and so dividing by 3 to get x gives \[x=\frac{\pi}{9}\]

OpenStudy (ybarrap):

2cos(3x) - 1 = 0 2cos(3x)=1 cos(3x) = 1/2 3x = acos(1/2) x = acos(1/2) / 3 = pi/3/3 = pi/6 Ans pi/6

OpenStudy (anonymous):

not quite, but close. last step was not right.

OpenStudy (ybarrap):

i.e pi/9

OpenStudy (anonymous):

look at the unit circle on the last page of the cheat sheet. find the places on the unit circle where the first coordinate is 1/2 then look at the angle. those will be the angles with cosine of 1/2

OpenStudy (anonymous):

Thanks @satelite73 soooooo much! You saved my life. I have another question if you don't mind